# SpeechJudge: Towards Human-Level Judgment for Speech Naturalness

### v1

@@ 5.1 Methodology
We develop SpeechJudge-GRM based on Qwen2.5-Omni-7B (Thinker) [56]. Inspired by the powerful capabilities of RL with the verifiable reward (RLVR) [42, 11], our natural initial approach is to treat the human preference $y_{\mathcal{H}}$ for the pair $(a_{1},a_{2})$ as a verifiable reward, and launch a RLVR training based on Qwen2.5-Omni. However, in practice, we find that the instruction-following reasoning capabilities of Qwen2.5-Omni are very weak (more detailed discussions can be found in Appendix 10). Therefore, we adopt a two-stage post-training process (“SFT + RL”) to develop SpeechJudge-GRM (Figure 4). We describe the details as follows.
@@
SFT Stage We consider SFT as a “cold start” stage to improve the Qwen2.5-Omni’s instruction-following, reasoning, and speech naturalness understanding capabilities. We select Gemini-2.5-Flash [8]—one of the leading closed-source models on SpeechJudge-Eval (Table 2)—to serve as a teacher model, and instruct it to generate the CoT data. Specifically, for each sample $d=(t,a_{1},a_{2},y_{\mathcal{H}})$ from SpeechJudge-Data, we use the CoT prompt from Table 1 (denoted as $\mathbf{I}_{CoT}$) to instruct Gemini-2.5-Flash to generate a rationale-based output (denoted as $\mathbf{O}_{teacher}$). We then extract the preference judgment ($y_{\mathcal{M}}$) from this output. For samples where Gemini-2.5-Flash’s preference is consistent with the human (i.e., $y_{\mathcal{M}}=y_{\mathcal{H}}$), we concatenate the CoT prompt and the model’s output, $\mathbf{I}_{CoT},\mathbf{O}_{teacher}]$, to create a data point for our SFT dataset. Conversely, we consider the sample $d$ a challenging case and reserve the prompt $\mathbf{I}_{CoT}$ for the second-stage RL dataset. During the SFT stage, for each training sample $\mathbf{I}_{CoT},\mathbf{O}_{teacher}]$, we perform the next token prediction only on the segment $\mathbf{O}_{teacher}$.
RL Stage We treat the annotated human preference as a verifiable reward, and, building on the SFT model, we further trained it using the GRPO algorithm [42]. Specifically, for each sample $d=(t,a_{1},a_{2},y_{\mathcal{H}})$ in the RL dataset, we adopt the CoT prompt to instruct the policy model to conduct multiple rollouts during each iteration. For the $i$-th rollout, we parse the model’s preference for $(a_{1},a_{2})$, denoted as $y_{\mathcal{M}}^{i}$. Following [34], we use an accuracy-based rule to calculate the reward: the reward is 1 if $y_{\mathcal{M}}^{i}=y_{\mathcal{H}}$, and -1 otherwise. In other words, during the RL stage, we only constrain the model’s final naturalness judgment to align with human preferences, allowing the model to autonomously optimize its reasoning and rationale generation capabilities.
We denote the training dataset of SpeechJudge-GRM as SpeechJudge-Data (train). Its construction process is as follows (see Appendix 7.2 for more details). Based on the raw SpeechJudge-Data, we first filter out all samples at the Full Disagreement (FD) level. For the other samples—at the FA, WA, and WD levels—we apply a majority voting principle among annotators to determine the final label for each. We then further exclude samples with a “Tie” label, using only the remaining preference data to form the SpeechJudge-Data (train). We use LoRA [20] to fine-tune the GRM during both the SFT and RL stages. Other experimental setup details are provided in Appendix 11.
@@ 5.2 Effectiveness of SpeechJudge-GRM on Naturalness Judgement
To verify the effectiveness of SpeechJudge-GRM for naturalness judgment, we evaluate it on the SpeechJudge-Eval benchmark. We develop SpeechJudge-BTRM as a baseline, which utilizes the BTRM paradigm [4, 39] by adding a linear layer on Qwen2.5-Omni-7B (Thinker) to produce a single scalar reward prediction. SpeechJudge-BTRM also uses LoRA fine-tuning and uses the same training data as SpeechJudge-GRM.
From the results of Table 3, we can observe that: (1) The SpeechJudge-BTRM achieves a 72.7% agreement with human preferences on SpeechJudge-Eval, a level of performance comparable to the initial development of BTRMs in the textual LLM RLHF field [44, 3, 37]. (2) After conducting SFT training with the CoT data, the accuracy of SpeechJudge-GRM (SFT) reaches 75.3%. Besides, further RLVR training improves the final model SpeechJudge-GRM (SFT+RL) to an accuracy of 77.2%. (3) Due to the generative nature of the GRM, we can further enhance the accuracy of SpeechJudge-GRM using inference-time scaling. For example, by using majority voting across 10 outputs instead of just one, the accuracy is improved by approximately 2 percentage points (75.3% $\rightarrow$ 77.6%; 77.2% $\rightarrow$ 79.4%). These results collectively verify the effectiveness of our proposed SpeechJudge-GRM for judging speech naturalness.
@@ 5.3 High-Quality Sample Selection based on SpeechJudge-GRM
We investigate the effect of SpeechJudge-based reward models for high-quality sample selection. We use the hard cases from SeedTTS-Eval [2] and the code-switching cases from Amphion-TTS-Eval [60] as target texts. For each text, we instruct the Qwen2.5-Omni-7B (Talker) [57] to generate 100 speeches. We then ask human subjects to compare the best-of-100 output—as selected by either SpeechJudge-BTRM or SpeechJudge-GRM—against a randomly sampled output. The evaluation measures the win/lose/tie ratios based on speech naturalness. From Figure 5, we observe that the best-of-100 samples selected by both SpeechJudge-BTRM and SpeechJudge-GRM are more likely to outperform a randomly selected sample from the same set. This finding demonstrates the advantage of using the SpeechJudge-Data corpus for training human-aligned reward model. Furthermore, SpeechJudge-GRM exhibits better performance than SpeechJudge-BTRM, which highlights the superiority of the proposed GRM.
@@ 5.4 Post-Training of Zero-Shot TTS based on SpeechJudge-GRM
We investigate the effect of using SpeechJudge-GRM as a reward function for post-training of TTS model. Specifically, we develop a new zero-shot TTS model, Qwen2.5-0.5B-TTS, to serve as the base model, which was not involved in the construction of the SpeechJudge-Data. This model is based on Qwen2.5-0.5B [57], adopts the classic two-stage “AR+Diffusion” architecture [2, 13], uses the speech tokenizer from DualCodec [31], and is pre-trained on the Emilia dataset [19].
@@
Based on this pre-trained model, we design four comparative methods: (1) w/ INTP: We use the intelligibility preference dataset, INTP [61], to perform offline DPO alignment [39]. (2) w/ SpeechJudge-Data: We use the SpeechJudge-Data (train) to perform offline DPO alignment. (3) w/ SpeechJudge-GRM (offline): We use SpeechJudge-GRM as an offline preference data annotator. We take all speech pairs from the INTP dataset and re-annotate their preference labels using SpeechJudge-GRM, then perform offline DPO alignment on the resulting data. (4) w/ SpeechJudge-GRM (online): We use SpeechJudge-GRM as a reward function for the online DPO algorithm [18]. The training data consists of only the prompts from INTP (i.e., the target texts and speech references for zero-shot TTS).
@@
We use SeedTTS-Eval [2] and Amphion-TTS-Eval [61, 60, 62] as evaluation sets. We present the objective results (WER and SIM) in Table 6 and the subjective results in Figure 6. We observe that both intelligibility and naturalness are enhanced for all the four methods after post-training. Additionally, the post-training method based on SpeechJudge-GRM achieves a greater improvement in naturalness (Figure 6(a)). Besides, the SpeechJudge-based methods could match or lead to a slight improvement in speaker similarity (Figure 6(b)).

### v2

@@ 5.1 Methodology
We develop SpeechJudge-GRM based on Qwen2.5-Omni-7B (Thinker) [qwen2.5-omni]. Inspired by the powerful capabilities of RL with the verifiable reward (RLVR) [grpo, deepseek-r1], our natural initial approach is to treat the human preference $y_{\mathcal{H}}$ for the pair $(a_{1},a_{2})$ as a verifiable reward, and launch a RLVR training based on Qwen2.5-Omni. However, in practice, we find that the instruction-following reasoning capabilities of Qwen2.5-Omni are very weak (more detailed discussions can be found in Appendix 11). Therefore, we adopt a two-stage post-training process (“SFT + RL”) to develop SpeechJudge-GRM (Figure 4). We describe the details as follows.
SFT Stage We consider SFT as a “cold start” stage to improve the Qwen2.5-Omni’s instruction-following, reasoning, and speech naturalness understanding capabilities. We select Gemini-2.5-Flash [gemini2.5]—one of the leading closed-source models on SpeechJudge-Eval (Table 2)—to serve as a teacher model, and instruct it to generate the CoT data. Specifically, for each sample $d=(t,a_{1},a_{2},y_{\mathcal{H}})$ from SpeechJudge-Data, we use the CoT prompt from Table 1 (denoted as $\mathbf{I}_{CoT}$) to instruct Gemini-2.5-Flash to generate a rationale-based output (denoted as $\mathbf{O}_{teacher}$). We then extract the preference judgment ($y_{\mathcal{M}}$) from this output. For samples where Gemini-2.5-Flash’s preference is consistent with the human (i.e., $y_{\mathcal{M}}=y_{\mathcal{H}}$), we concatenate the CoT prompt and the model’s output, $\mathbf{I}_{CoT},\mathbf{O}_{teacher}]$, to create a data point for our SFT dataset. Conversely, we consider the sample $d$ a challenging case and reserve the prompt $\mathbf{I}_{CoT}$ for the second-stage RL dataset. During the SFT stage, for each training sample $\mathbf{I}_{CoT},\mathbf{O}_{teacher}]$, we perform the next token prediction only on the segment $\mathbf{O}_{teacher}$.
RL Stage We treat the annotated human preference as a verifiable reward, and, building on the SFT model, we further trained it using the GRPO algorithm [grpo]. Specifically, for each sample $d=(t,a_{1},a_{2},y_{\mathcal{H}})$ in the RL dataset, we adopt the CoT prompt to instruct the policy model to conduct multiple rollouts during each iteration. For the $i$-th rollout, we parse the model’s preference for $(a_{1},a_{2})$, denoted as $y_{\mathcal{M}}^{i}$. Following [deepseek-grm], we use an accuracy-based rule to calculate the reward: the reward is 1 if $y_{\mathcal{M}}^{i}=y_{\mathcal{H}}$, and -1 otherwise. In other words, during the RL stage, we only constrain the model’s final naturalness judgment to align with human preferences, allowing the model to autonomously optimize its reasoning and rationale generation capabilities.
We denote the training dataset of SpeechJudge-GRM as SpeechJudge-Data (train). Its construction process is as follows (see Appendix 8.2 for more details). Based on the raw SpeechJudge-Data, we first filter out all samples at the Full Disagreement (FD) level. For the other samples—at the FA, WA, and WD levels—we apply a majority voting principle among annotators to determine the final label for each. We then further exclude samples with a “Tie” label, using only the remaining preference data to form the SpeechJudge-Data (train). We use LoRA [lora] to fine-tune the GRM during both the SFT and RL stages. Other experimental setup details are provided in Appendix 12.
@@ 5.2 Effectiveness of SpeechJudge-GRM on Naturalness Judgement
To verify the effectiveness of SpeechJudge-GRM for naturalness judgment, we evaluate it on the SpeechJudge-Eval benchmark. We develop SpeechJudge-BTRM as a baseline, which utilizes the BTRM paradigm [bt-rm, dpo] by adding a linear layer on Qwen2.5-Omni-7B (Thinker) to produce a single scalar reward prediction. SpeechJudge-BTRM also uses LoRA fine-tuning and uses the same training data as SpeechJudge-GRM.
From the results of Table 3, we can observe that: (1) The SpeechJudge-BTRM achieves a 72.7% agreement with human preferences on SpeechJudge-Eval, a level of performance comparable to the initial development of BTRMs in the textual LLM RLHF field [openai-rm, RLHF-anthoropic, instructgpt]. (2) After conducting SFT training with the CoT data, the accuracy of SpeechJudge-GRM (SFT) reaches 75.3%. Besides, further RLVR training improves the final model SpeechJudge-GRM (SFT+RL) to an accuracy of 77.2%. (3) Due to the generative nature of the GRM, we can further enhance the accuracy of SpeechJudge-GRM using inference-time scaling. For example, by using majority voting across 10 outputs instead of just one, the accuracy is improved by approximately 2 percentage points (75.3% $\rightarrow$ 77.6%; 77.2% $\rightarrow$ 79.4%). These results collectively verify the effectiveness of our proposed SpeechJudge-GRM for judging speech naturalness.
@@ 5.3 High-Quality Sample Selection based on SpeechJudge-GRM
We investigate the effect of SpeechJudge-based reward models for high-quality sample selection. We use the hard cases from SeedTTS-Eval [seedtts] and the code-switching cases from Amphion-TTS-Eval [amphion] as target texts. For each text, we instruct the Qwen2.5-Omni-7B (Talker) [qwen2.5] to generate 100 speeches. We then ask human subjects to compare the best-of-100 output—as selected by either SpeechJudge-BTRM or SpeechJudge-GRM—against a randomly sampled output. The evaluation measures the win/lose/tie ratios based on speech naturalness. From Figure 5, we observe that the best-of-100 samples selected by both SpeechJudge-BTRM and SpeechJudge-GRM are more likely to outperform a randomly selected sample from the same set. This finding demonstrates the advantage of using the SpeechJudge-Data corpus for training human-aligned reward model. Furthermore, SpeechJudge-GRM exhibits better performance than SpeechJudge-BTRM, which highlights the superiority of the proposed GRM.
@@ 5.4 Post-Training of Zero-Shot TTS based on SpeechJudge-GRM
We investigate the effect of using SpeechJudge-GRM as a reward function for post-training of TTS model. Specifically, we develop a new zero-shot TTS model, Qwen2.5-0.5B-TTS, to serve as the base model, which was not involved in the construction of the SpeechJudge-Data. This model is based on Qwen2.5-0.5B [qwen2.5], adopts the classic two-stage “AR+Diffusion” architecture [seedtts, cosyvoice], uses the speech tokenizer from DualCodec [dualcodec], and is pre-trained on the Emilia dataset [emilia-large].
@@
Based on this pre-trained model, we design four comparative methods: (1) w/ INTP: We use the intelligibility preference dataset, INTP [intp], to perform offline DPO alignment [dpo]. (2) w/ SpeechJudge-Data: We use the SpeechJudge-Data (train) to perform offline DPO alignment. (3) w/ SpeechJudge-GRM (offline): We use SpeechJudge-GRM as an offline preference data annotator. We take all speech pairs from the INTP dataset and re-annotate their preference labels using SpeechJudge-GRM, then perform offline DPO alignment on the resulting data. (4) w/ SpeechJudge-GRM (online): We use SpeechJudge-GRM as a reward function for the online DPO algorithm [online-dpo]. The training data consists of only the prompts from INTP (i.e., the target texts and speech references for zero-shot TTS).
@@
We use SeedTTS-Eval [seedtts] and Amphion-TTS-Eval [intp, amphion, vevo2] as evaluation sets. We present the objective results (WER and SIM) in Table 14 and the subjective results in Figure 6. We observe that both intelligibility and naturalness are enhanced for all the four methods after post-training. Additionally, the post-training method based on SpeechJudge-GRM achieves a greater improvement in naturalness (Figure 6(a)). Besides, the SpeechJudge-based methods could match or lead to a slight improvement in speaker similarity (Figure 6(b)).
6 Limitations and Future Work
While SpeechJudge-Data and SpeechJudge-GRM represent a step toward human-aligned speech naturalness judges, several limitations remain and open up directions for future work.
Scope of data and annotators. Our corpus is constructed entirely from synthetic TTS outputs in Chinese, English, and Chinese–English code-switching, and our annotators are professional raters in China (native Mandarin speakers with high but still L2 English proficiency). As shown in Appendix 9.2, inter-annotator agreement is noticeably higher on Chinese than on English and mixed subsets, indicating that the current dataset primarily reflects the preferences of Chinese and Chinese–English bilingual listeners, and is tailored to TTS-style read speech rather than spontaneous conversation. Extending SpeechJudge-Data to more languages, speaking styles, and listener populations (including native speakers of other languages and more diverse cultural backgrounds) is an important direction for building more universal naturalness judges.
Residual failure cases. Our error analysis in Appendix 13.3 shows that SpeechJudge-GRM’s remaining mistakes on SpeechJudge-Eval concentrate on some specific trade-offs, such as clean but robotic vs. slightly noisy but lively speech, prosody vs. articulation, and extreme expressive styles like very high-F0 emotional speech or whispers. In these regimes the model can over-weight cleanness, under-weight style-appropriate prosody, or become effectively indifferent when preference gaps are extremely small. Future work could incorporate explicit modeling of recording conditions (e.g., background noise), style-aware priors, and targeted augmentation of expressive and cross-lingual examples to better capture these nuanced preferences.
CoT quality and teacher bias. The CoT capability of SpeechJudge-GRM is bootstrapped from a proprietary teacher (Gemini-2.5-Flash) via an SFT stage. Although our analyses in Appendix 13.2 suggest that GRM’s CoT is largely self-consistent and moderately faithful, the reasoning style still inherits biases from the teacher, and we do not perform large-scale human verification of the intermediate explanations. An interesting direction is to involve humans more directly in assessing and curating CoT rationales—similar in spirit to the concurrent SQ-LLM work [sq-llm], which incorporates human involvement in CoT annotation—for example, by collecting human-written or human-edited analyses, or learning from explicit feedback on explanation quality, and to explore alternative, fully open-source teachers for bootstrapping.
From coarse-grained to fine-grained naturalness. In this work, naturalness is annotated and modeled at the utterance level: for each sentence, annotators choose which of two speeches is more natural, and SpeechJudge-GRM outputs a single decision per pair. However, in real-world speech, naturalness is often highly non-uniform within an utterance—some segments sound very natural while others contain local artifacts, disfluencies, or prosodic issues. Our current formulation does not explicitly localize such fine-grained phenomena. A promising future direction is to collect segment-level or time-aligned human feedback and to train reward models that can produce not only utterance-level judgments but also fine-grained scores or rationales over time.